# Thermodynamic uncertainty relations in superconducting junctions

### v1

@@ Abstract
Quantum conductors attached to metallic reservoirs have been demonstrated to overcome the thermodynamic uncertainty relation (TUR), a trade-off relation between the amount of dissipation and the absence of charge and heat current fluctuations. Here, we report large TUR violations when superconducting reservoirs replace metallic ones. The coexistence of different transport processes, namely (multiple) Andreev reflection, where electrons and their retro-reflected holes create Cooper pairs, in addition to the normal quasiparticle transport is identified as the source for such TUR breakdowns. The large TUR violation is a remarkable advantage for building low dissipative and highly stable quantum thermal machines.
@@
Introduction.— The implementation of superconducting materials has become a key ingredient to design better and more functional quantum thermal machines. They have been proposed to be efficient for cooling performances [1, 2, 3, 4, 5], as well as for pumping or generating power [6, 7]. Besides, superconductors can exhibit remarkable thermoelectric properties in the nonlinear transport regime when electron-hole symmetry is broken [8, 9, 10, 11].
@@
Another advantageous aspect when dealing with superconducting elements is that they overcome the fundamental bound imposed by the thermodynamic uncertainty relation (TUR) [12, 13, 14, 15, 16, 17, 18, 19, 5]. The TUR establishes a trade-off relation between precision or uncertainty of a current traversing the system and the associated entropy production or dissipation. The TUR is a nonequilibrium relation derived from the Markovian evolution of a system and sets a more restrictive bound for the entropy production in terms of the noise to current ratio. In the case of charge transport, TUR dictates $S/I^{2}\geq 2k_{\rm B}/\dot{\Sigma}$ with $k_{\rm B}$ the Boltzmann constant, $\dot{\Sigma}$ the entropy production, $I$ the average charge current and $S$ the corresponding current fluctuations.
@@
TUR violations have been reported in quantum systems attributed to either quantum coherence or the breakdown of the local detailed balance [20, 21, 22, 5, 23, 24, 25]. Large deviations from the TUR are desirable for the quantum-enhanced performance of the engine operation, i.e., less dissipative and more stable machines [26, 19, 21, 27]. Besides, quite recently there has been a notable effort towards finding quantum systems exhibiting a prominent breakdown of the TUR [5, 24, 28, 23, 24]. Thus, for instance, a chain of quantum dots with a rectangular transmission was proposed to violate the standard TURs [28]. Additionally, moderate deviations in the TUR relation have been reported in hybrid normal-superconductor quantum dots in the weak tunnel coupling regime [5]. However, regimes where quantum coherence cannot be treated perturbatively have not been addressed yet when superconducting elements are present. Here, we investigate different superconducting devices where the TUR is largely violated, namely (i) hybrid setups in which a quantum conductor is coupled to a normal and a superconducting reservoir and (ii) quantum conductors coupled to two superconductors.
In this work, we demonstrate strong violations of the TUR due to the coexistence of different types of transport processes in normal-superconductor (NS), and superconductor-superconductor (SS) contacts. These processes correspond to single-quasiparticle tunneling (QP) and Andreev reflection (AR) or multiple Andreev reflection (MAR). Besides, our results show that the highest TUR departures arise for quasi-ballistic conductors in which the transmission is close to the unity without an energy dependence as occurs in a quantum point contact (QPC).
@@
Quantifying TUR deviations.— We follow Ref. [19] to find departures of the TUR. In the following, we assume $e=1$. For a generic isothermal two terminal device where there is a bias voltage $V$, entropy production is due to the Joule heating $\dot{\Sigma}=IV/T$ where $I$ is the charge current, and $T$ the contact temperature which is the same in both reservoirs. Deviations for the TUR, i.e., $\frac{S}{I^{2}}\dot{\Sigma}\geq 2k_{B}$ with $S$ the current fluctuations, are quantified by the TUR-breaking coefficient
@@
where $F^{*}=S/I$ is the Fano factor. The TUR-breaking coefficient $\mathcal{F}$ is negative for broken and positive (or zero) for non-broken TUR, where $\mathcal{F}=0$ corresponds to the classical limit. Current $I$ and noise $S$ for the charge can be expanded as
@@
$S$ $=$ $S_{0}+S_{1}V+\frac{1}{2!}S_{2}V^{2}+\mathcal{O}(V^{3}),\quad\quad$ (3)
in terms of the voltage. Here, $G_{1},G_{2},G_{3}$ are linear and nonlinear conductances and $S_{1},S_{2},S_{3}$ are the equilibrium and nonequilibrium noise terms. Using the nonlinear fluctuation relations [29, 30, 20, 31] together with the Johnson-Nyquist relation for the equilibrium noise (i.e., $S_{0}=2k_{\rm B}TG_{1}$) we obtain $\mathcal{F}=Vk_{\rm B}T/(G_{1})C_{\rm neq}+\mathcal{O}(V^{2})$ where the coefficient [19]
@@
quantifies TUR deviations. Namely when $C_{\text{neq}}<0$ (in the limit of small voltages), the TUR is violated.
@@
The stochastic nature of charge transfer in mesoscopic junctions allows us to express the current in an arbitrary mesoscopic junction as follows [32, 33]
$I=\int_{-\infty}^{\infty}\frac{dE}{h}\left(\sum_{n}np_{n}\right),$ (5)
@@
where $p_{n}=p_{n}(E,V)$ are charge-resolved probabilities of a transport process transferring $n$ charges across the junction that depend on the energy $E$ and on the voltage $V$. Current-current fluctuations are given by [32, 33]
$S=\int_{-\infty}^{\infty}\frac{dE}{h}\left[\sum_{n}n^{2}p_{n}-\left(\sum_{n}np_{n}\right)^{2}\right].$ (6)
We want to stress that the formulas for the current and noise fluctuations are completely general and also hold for junctions containing superconductors. To get further insight, we expand the charge-resolved probabilities in the applied voltage
$p_{n}=p_{n}^{0}+p_{n}^{1}V+p_{n}^{2}V^{2}/2+p_{n}^{3}V^{3}/6+\mathcal{O}(V^{4}).$ (7)
@@
By using the detailed balance condition $p_{-n}=p_{n}e^{-n\beta V}$ [30] where $\beta=1/(k_{\rm B}T)$ and inserting the expanded probabilities [see Eq. (7)] in the current and noise formula [see Eqs. (5), and (6)] we obtain the following result for the coefficient $C_{\rm neq}$ in Eq. (4)
$\frac{6C_{\rm neq}}{\beta^{3}}=\int\frac{dE}{h}\Bigg{(}$ $\sum_{n>0}n^{4}p_{n}^{0}(1-6p_{n}^{0})$
$-\sum_{m>n>0}12n^{2}m^{2}p_{n}^{0}p_{m}^{0}\Bigg{)},$ (8)
@@
which notably only depends on the zeroth order approximation of the transmission probabilities $p_{n}^{0}$.
@@
To exemplify Eq. (Thermodynamic uncertainty relations in superconducting junctions) in the context of superconducting junctions, we now consider a single-channel NS contact. The superconducting lead is assumed to be a conventional BCS superconductor with a temperature-dependent gap $\Delta\equiv\Delta(T)$. In a NS junction, there are two transport processes as illustrated in Fig. 1(a,b). For zero temperature, QPs can only be transferred for energies above the gap ($V\geq\Delta$) and transfer one charge which is shown in panel Fig. 1(a). In the subgap region ($V\leq\Delta$) transport can take place via Andreev reflection where two charges are transferred in one instance, see panel Fig. 1(b). For non-zero temperatures, QP transfer can also occur for in-gap voltages because of the thermal broadening of the corresponding Fermi functions. The two transport processes are associated with different transmission probabilities, namely $p_{\pm 1}$ corresponds to QP tunneling and $p_{\pm 2}$ to AR and all other probabilities are zero [34]. With this information at hand, the coefficient $C_{\rm neq}$ reads:
$\frac{6C_{\rm neq}^{\rm NS}}{\beta^{3}}=\int\frac{dE}{h}\left[p_{1}^{0}(1-6p_{1}^{0})+2^{4}p_{2}^{0}(1-6p_{2}^{0})-48p_{1}^{0}p_{2}^{0}\right].$ (9)
We see the QP part $p_{1}^{0}(1-6p_{1}^{0})$ which agrees with the result in Ref. [19] when $p_{2}^{0}=0$, i.e., in the absence of AR (or similar) processes. In addition, we have a new term that has the same form but is related to the Andreev reflection $2^{4}p_{2}^{0}(1-6p_{2}^{0})$. Furthermore, we get a term with a negative sign containing a product of both probabilities, namely $-48p_{1}^{0}p_{2}^{0}$, which generally increases the chance to break TUR. More importantly, the contribution of higher order processes [$n^{4}p_{n}^{0}(1-6p_{n}^{0})$] scales to the power of four in the charge, resulting in overall larger violations of TUR.
@@
In the case of SS contacts, new transport processes in the form of multiple Andreev reflection (MAR) take place. They correspond to probabilities $p_{\pm n}$ for $n>2$, which contribute to the current and noise (see Ref. [35, 36]) and therefore appear in Eq. (Thermodynamic uncertainty relations in superconducting junctions). The lowest order transport processes in an SS junction are depicted in Fig. 1(c), (d), (e) and (f) and correspond to QP ($p_{\pm 1}$), AR ($p_{\pm 2}$), MAR ($p_{\pm 3}$) and 4th order MAR ($p_{\pm 4}$) transferring one, two, three and four charges respectively.
@@
Figure 1: (a) Single-QP tunneling in which a QP tunnels into the empty density of states of the superconductor transferring one charge. The applied voltage has to exceed the SC gap, $V\geq\Delta$. (b) AR in which an electron is reflected as a hole inside the superconducting gap transferring two charges. The voltage is arbitrary, $V\geq 0$. (c) QP tunneling between two SCs at voltages above $2\Delta$. (d) AR from a voltage biased superconductor to another superconductor transferring two charges. The voltage has to exceed the SC gap, $V\geq\Delta$. (e) First MAR, where an incoming electron is reflected as a hole and that hole is retro-reflected as another electron, transferring a total of three charges. The voltage has to exceed $V\geq 2\Delta/3$. (f) Higher order MAR involves a process that transfers four charges with onset voltage $V\geq 2\Delta/4$. The voltage thresholds correspond to the case of zero temperature.
@@
QPC of constant transmission.— In the following we focus on single-channel NS and SS contacts characterized by an energy-independent normal state transmission coefficient $\tau$. These junctions have been realized with the help of superconducting atomic-size contacts in which both the current and the noise have been thoroughly studied [37, 38, 39, 40, 41, 42]. We make use of the framework of full counting statistics to obtain current, noise and charge resolved probabilities using Ref. [43] to compute the full result for the TUR-breaking coefficient in Eq. (1).
@@
Figure 2: (a) Critical transmission $\tau_{c}$ for breaking TUR for a NS junction as a function of the bias voltage $V/\Delta$ and the temperature $k_{\rm B}T/\Delta$. For areas with no color, TUR is not broken. (b) The minimal TUR breaking coefficient $\mathcal{F}_{\rm min}$ as a function of the dimensionless temperature $k_{\rm B}T/\Delta$ and the transmission $\tau$ for a NS junction. A negative coefficient indicates broken TUR. The black solid line indicates the phase boundary. The dotted line indicates the approximate phase boundary from the coefficient $C_{\rm neq}$ [see Eq. (9)]. (c) Same as in (a) for an SS junction. Dotted lines indicate special voltages of onsets of different transport processes, namely $V/\Delta=1,2/3,1/2$ for AR, MAR and 4th order MAR. Grey area shows a parameter realm that was not analyzed. (d) Same as in (b) for an SS junction with the phase boundary as a solid line.
Our purpose is to investigate whether TUR can be broken in a NS and SS system and portray the “phase” diagram of the system. We assign the phase of unbroken TUR if $\mathcal{F}\geq 0$ and the second phase of broken TUR if $\mathcal{F}<0$. To illustrate the results, we fix the voltage and temperature and vary the transmission $\tau$ of the QPC. For some value $\tau_{\rm c}$, which we denote as the critical transmission, the TUR inequality will be minimally broken ($\mathcal{F}<0$). It holds that for any $\tau<\tau_{\rm c}$, TUR is fulfilled whereas it is violated for $\tau>\tau_{\rm c}$.
@@
In Fig. 2(a), we show the critical transmission $\tau_{\rm c}$ as a function of the dimensionless temperature $k_{\rm B}T/\Delta$ and voltage $V/\Delta$ for the NS case. Notice that the TUR only becomes broken upon really high transmission values ($\tau_{c}\approx 0.91$). The higher the temperature, the smaller the critical transmission which also weakly depends on the voltage. We want to stress that no energy dependence of the transmission is needed for breaking TUR as the source of nonlinearity needed to violate TUR arises from the coexistence of QP tunneling and AR in addition to the non-constant density of states of the SC. This is in contrast to the results in Ref. [19], where some kind of nonlinearity of the transmission is needed. For low temperatures, TUR is only broken for $V\approx\Delta$ and only for $\tau\approx 1$, which suggests that the breaking of TUR is rooted in the coexistence of the QP and AR, which is the main source of nonlinearity. In particular, we find that the TUR is broken at the onset of QP tunneling for arbitrarily low temperatures. In particular, at that special voltage, the TUR-breaking coefficient scales with a power-law $\mathcal{F}(V=\Delta)=-(k_{\rm B}T/\Delta)^{3/2}$ (for more details see Ref. [44]).
@@
To quantify the violation of TUR, we consider the voltage at which TUR is maximally broken and define the minimal TUR-breaking coefficient $\mathcal{F}_{\rm min}\equiv{\rm min}_{V}\mathcal{F}$, which is positive (or zero) for unbroken TUR and negative for broken TUR. In addition, the more negative $\mathcal{F}_{\rm min}$ the more TUR is broken. In Fig. 2(b), this coefficient is plotted as a function of the transmission and the temperature. The solid line indicates the phase boundary between the unbroken and broken TUR phase. Notice that the TUR is more broken the higher the transmission is. To understand the reason why TUR is largely broken in the case of a NS junction we consider the approximate measure for breaking TUR, namely $C_{\rm neq}^{\rm NS}$ in Eq. (9). In the case of a NS junction, the transmission probabilities $p_{1,2}$ are analytical (see Ref. [43]) and $C_{\rm neq}^{\rm NS}$ can be easily calculated. The phase boundary arising from this approximation is shown as a dotted line in panel Fig. 2(b). It can be seen that there is a good correspondence in a qualitative sense for higher temperatures, whereas there are deviations for lower ones. The reason behind this is that in the case of larger temperatures, the temperature smearing is responsible for the form of the transmission probabilities even at finite bias. When the temperature is large, the form of $p_{1/2}^{0}$ is similar to $p_{1/2}(V>0)$ resulting in the approximation in Eq. (Thermodynamic uncertainty relations in superconducting junctions) to be valid for larger voltages (for more details see Ref. [44]). The coefficients $n^{4}p_{n}(1-6p_{n})$ scale quartically with the charge which renders a stronger violation of the TUR for highly transmissive AR processes coexisting with low transmissive QP processes. Furthermore, an additional yet small negative contribution of $C_{\rm neq}^{\rm NS}$ comes from the coupling term $\propto-p_{1}^{0}p_{2}^{0}$.
@@
For completeness we have addressed in Ref. [44] the case of a quantum conductor that exhibits a highly energy dependent transmission coefficient. In particular, we have studied the case of a quantum dot where the transmission coefficient corresponds to a Breit-Wigner resonance centered in the quantum dot level $\epsilon_{0}$ with a tunneling broadening given by $\Gamma$, i.e., $\tau(E)=\Gamma^{2}/[(E-\epsilon_{0})^{2}+\Gamma^{2}]$. However, we find that this added nonlinearity does not further break TUR, see Ref. [44].

### v2

@@ Abstract
It has been recently predicted that the thermodynamic uncertainty relation (TUR), which expresses a trade-off between the amount of dissipation and the absence of fluctuations, can be violated in the context of charge transport in normal quantum conductors. However, the required conditions are difficult to meet and no experimental violation has been yet reported in mesoscopic quantum transport experiments. Here, we predict that TUR can be largely violated in superconducting contacts without introducing an energy dependency in the normal transmission coefficient. This violation originates from the coexistence of tunneling processes involving different numbers of transmitted charges, namely quasiparticle tunneling and Andreev reflections. The superconducting junctions invoked here have been fabricated by many groups and constitute an ideal platform to demonstrate TUR’s breaking in charge transport experiments.
@@
Introduction.— The implementation of superconducting materials has become a key ingredient to design better and more functional quantum thermal machines. They have been proposed to be efficient for cooling performances [nahum1994electronic, leivo1996efficient, PhysRevB.98.241414, PhysRevB.107.245412, PhysRevResearch.5.043041], as well as for pumping or generating power [PhysRevB.106.115419, PhysRevB.103.235434]. Besides, superconductors can exhibit remarkable thermoelectric properties in the nonlinear transport regime when electron-hole symmetry is broken [PhysRevLett.124.106801, germanese2022bipolar, PhysRevResearch.6.L012022, PhysRevB.106.245413].
@@
Another advantageous aspect when dealing with superconducting elements is that they overcome the fundamental bound imposed by the thermodynamic uncertainty relation (TUR) [Barato2015, Todd2016, Horowitz2020, Pietz2016, Pietzonka16b, Pietz2018, Timp2019, PhysRevB.98.155438, PhysRevResearch.5.043041]. The TUR establishes a trade-off relation between uncertainty of a current traversing the system and the associated entropy production or dissipation. The TUR is a nonequilibrium relation derived from the Markovian evolution of a system and sets a more restrictive bound for the entropy production in terms of the noise to current ratio. In the case of charge transport, TUR dictates $S/I^{2}\geq 2k_{\rm B}/\dot{\Sigma}$ with $k_{\rm B}$ the Boltzmann constant, $\dot{\Sigma}$ the entropy production, $I$ the average charge current and SS the corresponding current fluctuations.
@@
TUR violations have been predicted in quantum systems attributed to either quantum coherence or the breakdown of the local detailed balance [Andrieux_2007, PhysRevLett.120.090601, PhysRevX.11.021013, PhysRevResearch.5.043041, PhysRevResearch.5.013038, PhysRevResearch.5.023155, PhysRevB.28.1655]. Large deviations from the TUR are desirable for the quantum-enhanced performance of the engine operation, i.e., less dissipative and more stable machines [PhysRevB.98.085425, PhysRevB.98.155438, PhysRevLett.120.090601, PhysRevE.99.062141]. Besides, quite recently there has been a notable effort towards finding quantum systems exhibiting a prominent breakdown of the TUR [PhysRevResearch.5.043041, PhysRevResearch.5.023155, PhysRevB.104.045424, PhysRevResearch.5.013038, PhysRevResearch.5.023155, PhysRevB.108.115422]. Thus, for instance, a chain of quantum dots with a rectangular transmission was proposed to violate the standard TURs [PhysRevB.104.045424]. Additionally, moderate deviations in the TUR relation have been predicted in hybrid normal-superconductor quantum dots in the weak tunnel coupling regime [PhysRevResearch.5.043041]. However, all these proposals require stringent conditions and all attempts to observe a violation of TUR in quantum conductors have failed thus far, see e.g. Ref. [PhysRevB.101.195423].
Here, we predict that TUR can be largely violated in few-channel quantum point contacts involving superconducting leads, which have already been realized e.g. with the help of superconducting atomic-size contacts. We show that the strong violations of TUR in these junctions are due to the coexistence of different transport processes, namely single-quasiparticle tunneling (QP) and (multiple) Andreev reflections (AR). We find that the largest TUR violations occur for highly transmissive contacts and no energy dependence of the normal state transmission is required, contrary to what happens in normal quantum point contacts [PhysRevB.98.085425, PhysRevB.104.045424].
@@
Quantifying TUR deviations.— In what follows we present a general discussion to explain what makes superconductors special in the context of TUR. For this purpose, we consider a generic isothermal two-terminal quantum device that may contain normal metals (N) and/or superconductors (S). If we apply a bias voltage $V$, entropy production due to the Joule heating is given by $\dot{\Sigma}=IV/T$, where $I$ is the charge current and $T$ the temperature of the electrodes. Deviations for the TUR, i.e., $\frac{S}{I^{2}}\dot{\Sigma}\geq 2k_{B}$ with $S$ being the current fluctuations, can be quantified by the TUR-breaking coefficient
@@
where $F^{*}=S/I$ is the Fano factor and we have assumed that the electron charge is unity. The coefficient $\mathcal{F}$ is negative for broken TUR and positive (or zero) for unbroken TUR, where $\mathcal{F}=0$ corresponds to the classical limit. Below, we shall present a quantitative analysis of $\mathcal{F}$ valid for arbitrary range of parameters, but to get an insight into the conditions required to break TUR, we now follow Ref. [PhysRevB.98.085425] and focus on the low-bias regime in which it is possible to make analytical progress. With this idea in mind, we make a Taylor expansion in the bias of the current and noise
@@
$S$ $=$ $S_{0}+S_{1}V+\frac{1}{2!}S_{2}V^{2}+\mathcal{O}(V^{3}).\quad\quad$ (3)
Here, $G_{1},G_{2},G_{3}$ are linear and nonlinear conductances and $S_{0},S_{1},S_{2}$ are the equilibrium and nonequilibrium noise terms. Using the nonlinear fluctuation relations [PhysRevLett.101.136805, PhysRevB.72.235328, Andrieux_2007, PhysRevLett.108.246603] together with the Johnson-Nyquist relation for the equilibrium noise (i.e., $S_{0}=2k_{\rm B}TG_{1}$), we obtain $\mathcal{F}=Vk_{\rm B}T/(G_{1})C_{\rm neq}+\mathcal{O}(V^{2})$ where the coefficient [PhysRevB.98.155438]
@@
quantifies TUR deviations. Namely when $C_{\text{neq}}<0$ (in the limit of small voltages), TUR is violated.
@@
The stochastic nature of charge transfer in mesoscopic junctions allows us to express the current in an arbitrary mesoscopic junction as follows [Levitov1996, Klich2003]
$I=\int\frac{dE}{h}\left(\sum_{n}np_{n}\right),$ (5)
@@
where $p_{n}=p_{n}(E,V,T)$ are charge-resolved probabilities of a transport process transferring $n$ charges across the junction that depends on the energy $E$, voltage $V$ and temperature $T$. The corresponding current fluctuations are given by [Levitov1996, Klich2003]
$S=\int\frac{dE}{h}\left[\sum_{n}n^{2}p_{n}-\left(\sum_{n}np_{n}\right)^{2}\right].$ (6)
These formulas for $I$ and SS hold for junctions containing normal metals and/or superconductors. For normal contacts, the transport is solely due to single-electron tunneling ($|n|=1$), while with S electrodes multiple electron ($|n|>1$) tunneling processes can coexist. This is what makes superconductors special, as we are about to show.
To get further insight, we expand the charge-resolved probabilities in the applied voltage
$p_{n}=p_{n}^{(0)}+p_{n}^{(1)}V+p_{n}^{(2)}V^{2}/2+p_{n}^{(3)}V^{3}/6+\mathcal{O}(V^{4}).$ (7)
@@
By using the detailed balance condition $p_{-n}=p_{n}e^{-n\beta V}$ [PhysRevB.72.235328] where $\beta=1/(k_{\rm B}T)$ and inserting the expanded probabilities [see Eq. (7)] in the current and noise formula [see Eqs. (5), and (6)] we obtain the following result for the coefficient $C_{\rm neq}$ in Eq. (4)
$\frac{6C_{\rm neq}}{\beta^{3}}=\int\frac{dE}{h}\Bigg($ $\sum_{n>0}n^{4}p_{n}^{(0)}(1-6p_{n}^{(0)})$
$-\sum_{m>n>0}12n^{2}m^{2}p_{n}^{(0)}p_{m}^{(0)}\Bigg),$ (8)
@@
which notably only depends on the zeroth order approximation of the transmission probabilities $p_{n}^{(0)}$.
@@
To exemplify Eq. (Thermodynamic uncertainty relations in superconducting junctions) in the context of superconducting junctions, we now consider a single-channel NS contact. The superconducting lead is assumed to be a conventional BCS superconductor with a temperature-dependent gap $\Delta\equiv\Delta(T)$. In this case, there are two transport processes as illustrated in Fig. 1(a,b). For zero temperature, QPs can only be transferred for energies above the gap ($V\geq\Delta$) and transfer one charge which is shown in panel Fig. 1(a). In the subgap region ($V\leq\Delta$) transport can take place via Andreev reflection where two charges are transferred in one instance, see panel Fig. 1(b). For non-zero temperatures, QP transfer can also occur for in-gap voltages because of the thermal broadening of the corresponding Fermi functions. The two processes are associated with different transmission probabilities, namely $p_{\pm 1}$ corresponds to QP tunneling and $p_{\pm 2}$ to AR and all other probabilities are zero [Belzig2003]. With this information at hand, the coefficient $C_{\rm neq}$ reads:
$\frac{6C_{\rm neq}^{\rm NS}}{\beta^{3}}{=}\int\frac{dE}{h}\left[p_{1}^{(0)}(1{-}6p_{1}^{(0)}){+}2^{4}p_{2}^{(0)}(1{-}6p_{2}^{(0)}){-}48p_{1}^{(0)}p_{2}^{(0)}\right].$ (9)
This expression reduces to the result in Ref. [PhysRevB.98.155438] for normal contacts ($p_{2}^{(0)}=0$). The new contributions in the NS case, second and third terms, are related to the AR and the coexistence of the two processes, respectively. The sign of $C_{\rm neq}^{\rm NS}$ depends on the precise energy dependence of $p^{(0)}_{1,2}$. One can gain some intuition by analyzing an NS contact with an energy-independent normal transmission coefficient. In this case the $p^{(0)}_{1,2}$ coefficients can be computed e.g. with the well-known BTK model [PhysRevB.25.4515]. As we show in Ref. [SM], such an analysis suggests that two requirements must be simultaneously met to significantly break TUR: (i) coexistence of the two processes at a given bias and temperature and (ii) high transmissions. The coexistence of the two processes is responsible for a large negative contribution of the cross term $48p_{1}^{(0)}p_{2}^{(0)}$ in Eq. (9), while high transmissions are necessary to have high probabilities and make the terms $n^{4}p_{n}^{(0)}(1-6p_{n}^{(0)})$ negative. The dependence of the zero-bias probabilities on transmission and temperature, as well as a discussion on the role of the different terms in Eq. (9), can be found in Ref. [SM]. This intuition will be further confirmed below by our numerically exact results at arbitrary bias and we shall use those results to test the validity of Eq. (9).
@@
In the case of SS contacts, new transport processes in the form of multiple Andreev reflections (MAR) can take place. Their corresponding probabilities $p_{\pm n}$ for $n>2$ contribute to the current and noise (see Ref. [Cuevas2003, Cuevas2004]) and therefore appear in Eq. (Thermodynamic uncertainty relations in superconducting junctions). The lowest order transport processes in an SS junction are depicted in Fig. 1(c-f) and correspond to QP ($p_{\pm 1}$), AR ($p_{\pm 2}$), MAR ($p_{\pm 3}$) and 4th order MAR ($p_{\pm 4}$) transferring one, two, three and four electrons respectively.
@@
Figure 1: (a) Single-QP tunneling in which a QP tunnels into the empty density of states of the superconductor transferring one electron. The applied voltage has to exceed the SC gap, $V\geq\Delta$. (b) AR in which an electron is reflected as a hole inside the superconducting gap transferring two electrons. The voltage is arbitrary, $V\geq 0$. (c) QP tunneling between two SCs at voltages above $2\Delta$. (d) AR from a voltage biased superconductor to another superconductor transferring two electrons. The voltage has to exceed the SC gap, $V\geq\Delta$. (e) First MAR, where an incoming electron is reflected as a hole and that hole is retro-reflected as another electron, transferring a total of three electrons. The voltage has to exceed $V\geq 2\Delta/3$. (f) Higher order MAR involves a process that transfers four electrons with onset voltage $V\geq 2\Delta/4$. The voltage thresholds correspond to the case of zero temperature.
@@
QPC of constant transmission.— In the following we focus on single-channel NS and SS contacts characterized by an energy-independent normal state transmission coefficient $\tau$. These junctions have been realized with the help of superconducting atomic-size contacts in which both the current and the noise have been thoroughly studied [Scheer1997, Scheer1998, Cuevas1998, Cron2001, Senkpiel2020, Senkpiel2020a]. We make use of the framework of full counting statistics to obtain current, noise and charge resolved probabilities using Ref. [Ohnmacht2023] to compute the full result for the TUR-breaking coefficient in Eq. (1).
@@
Figure 2: (a) Critical transmission $\tau_{c}$ for breaking TUR for a NS junction as a function of the bias voltage $V/\Delta$ and the temperature $k_{\rm B}T/\Delta$. For areas with no color, TUR is not broken. (b) The minimal TUR breaking coefficient $\mathcal{F}_{\rm min}$ as a function of the dimensionless temperature $k_{\rm B}T/\Delta$ and the transmission $\tau$ for a NS junction. A negative coefficient indicates broken TUR. The black solid line indicates the phase boundary. The dotted line indicates the approximate phase boundary from the coefficient $C_{\rm neq}$ [see Eq. (9)]. The inset shows a zoom-in into the region where TUR is broken. (c) Same as in (a) for an SS junction. Dotted lines indicate special voltages of onsets of different transport processes, namely $V/\Delta=1,2/3,1/2$ for AR, MAR and 4th order MAR. Grey area shows a parameter realm that was not analyzed. (d) Same as in (b) for an SS junction with the phase boundary as a solid line. The inset shows a zoom-in into the region where TUR is broken.
@@
Our purpose is to investigate whether TUR can be broken in a NS and SS system and portray the “phase” diagram of the system. We assign the phase of unbroken TUR if $\mathcal{F}\geq 0$ and the second phase of broken TUR if $\mathcal{F}<0$. To illustrate the results, we fix the voltage and temperature and vary the transmission $\tau$ of the QPC. We define $\tau_{\rm c}$ as the critical transmission for which the lower bound of TUR inequality is satisfied ($\mathcal{F}=0$). It holds that for any $\tau<\tau_{\rm c}$, TUR is fulfilled whereas it is violated for $\tau>\tau_{\rm c}$. In Fig. 2(a) we show $\tau_{\rm c}$ as a function of the dimensionless temperature $k_{\rm B}T/\Delta$ and voltage $V/\Delta$ for the NS case. Notice that the TUR only breaks at high transmissions ($\tau_{c}\approx 0.91$), as anticipated in the discussion of Eq. 9. Note also that $\tau_{\rm c}$ decreases upon increasing the temperature, which is due to the larger overlap between the two process induced by the finite temperature. We want to stress that no energy dependence of the transmission is needed for breaking TUR. The source of nonlinearity of the transport characteristics ($I$ and $S$) as a function of the bias that is required to violate TUR arises from the coexistence of QP tunneling and AR in addition to the non-constant density of states of the SC. This is in contrast to the results in Ref. [PhysRevB.98.155438], where an energy dependence of the transmission is needed. For low temperatures, TUR is only broken for $V\approx\Delta$ and only for $\tau\approx 1$. This indicates that the breaking of TUR is rooted in the coexistence of the QP and AR (for more details see Ref. [SM]).
@@
To quantify the violation of TUR, we consider the voltage at which TUR is maximally broken and define the minimal TUR-breaking coefficient $\mathcal{F}_{\rm min}\equiv{\rm min}_{V}\mathcal{F}$, which is positive (or zero) for unbroken TUR and negative for broken TUR. In addition, the more negative $\mathcal{F}_{\rm min}$ the more TUR is broken. In Fig. 2(b), this coefficient is plotted as a function of the transmission and the temperature. The solid line indicates the phase boundary between the unbroken and broken TUR phase. Notice that the TUR is more broken the higher the transmission is. These findings can be understood with the help of the low-bias analysis above, see Eq. (9). The probabilities $p_{1,2}$ are analytical in the NS (see Ref. [Ohnmacht2023, SM]) and $C_{\rm neq}^{\rm NS}$ can be easily calculated. The phase boundary obtained with this approximation is shown as a dotted line in panel Fig. 2(b). It qualitatively reproduces the exact results, especially at high temperatures where the energy dependence of $p_{1,2}$ is mainly determined by the smearing of the Fermi functions (for more details see Ref. [SM]). On the other hand, the fact that high transmissions are required to have a stronger violation of TUR can be understood from the contribution $2^{4}p^{(0)}_{2}(1-6p^{(0)}_{2})$ in Eq. (9). This term scales quartically with the charge, which renders a stronger violation for highly probable AR processes coexisting with low probability QP processes. Furthermore, an important contribution to $C_{\rm neq}^{\rm NS}$ comes also from the cross term $48p_{1}^{(0)}p_{2}^{(0)}$.
@@
For completeness, we also address in Ref. [SM] the case of an NS contact featuring an energy dependent transmission coefficient. In particular, we have studied the case in which the leads are coupled via a quantum dot. The transmission coefficient in this case corresponds to a Breit-Wigner resonance centered in the energy of the dot level $\epsilon_{0}$ with a tunneling broadening given by $\Gamma$, i.e., $\tau(E)=\Gamma^{2}/[(E-\epsilon_{0})^{2}+\Gamma^{2}]$. However, we find that this added energy dependence does not further break TUR, see Ref. [SM], which we attribute to the fact that the resonant level tends to reduce the high effective transmission at some energies that is needed to break TUR.